A hockey puck lies still on ice. Another hockey puck is hit and is on course of collision with it with an initial velocity of +v. The pucks have the same mass and have glue around their edges. Neglect friction with ice. Predict the velocity that the puck which was initially at rest will have after collision, knowing that the pucks will stick together:

a) –v

b) +v

c) –v/2

d) +v/2




1
Expert's answer
2022-04-11T12:03:49-0400

"u=+\frac{mv}{m+m}=+\frac v2."